Raul can drive 3 times as fast as Jose can ride his bicycle. If it takes Jose 2 hours longer than Raul to travel 24 miles, hoe fast can Jose ride his bicycle?
Answer by mananth(16949)

let jose ride at x mph
Time taken by jose for 24 miles = 24/x
raul rides at 3x mph
time taken by raul for 24 miles = 24/3x
24/x-24/3x=2
72-24 = 3x*2
48 = 6x
x= 8 mph which is jose's speed